OpenStudy (dmezzullo):

Evaluate: [(51 + 3) - 32] ÷ 9 · 2 52 13.3 2.5 10

OpenStudy (anonymous):

22 dived by 9 x 2

OpenStudy (anonymous):

2.44444444444444444x 2=4.8-reapeated9

OpenStudy (anonymous):

22 / 9 * 2

OpenStudy (anonymous):

so 51+3 = 54-32=22/9*2

OpenStudy (anonymous):

Yup